1/5t-0.4+2/5t=0.6-1/10t also can u show me how to check it please and thanks
Answer by stanbon(75887)

(1/5)t-0.4+(2/5)t = 0.6-(1/10)t
--------------
(3/5)t = 1 - (1/10)t
(6/10)t = 1 - (1/10)t
(7/10)t = 1
7t = 10
t = 10/7
------------------
(1/5)t-0.4+(2/5)t = 0.6-(1/10)t
Checking the answer:
(1/5)(10/7)-0.4+(2/5)(10/7) = 0.6-(1/10)(10/7)
(2/7) - 0.4 + (4/7) = 0.6 - (1/7)
(7/7) = 1
1 = 1
===============
Cheers,
Stan H.